# Compelling ReLU Networks to Exhibit Exponentially Many Linear Regions at Initialization and During Training

### v6

@@ A.9 Learning Rates
Figure 13: Learning $y=x^{3}$ at various learning rates for 1000 epochs. The figure shows the log losses of the best model of 30 for networks trained according to the methods in this paper versus random initialization and gradient descent (“Default Network”). For high learning rates, neither learns because the steps are too large. For low learning rates (below $10^{-5}$), the steps are too small, and our networks are likely deriving an implicit advantage from being forcibly initialized to a convex function. Both methods are able to converge for learning rates in between $10^{-4}$ and $10^{-2}$; one could run for more epochs to see a similar advantage of our method for smaller learning rates, if desired.
@@
A.10 Real-World Data and Classification Problems
@@
Nonetheless, while we believe there is much room for future work to improve higher-dimensional results, Figures 14 and 15 already show impressive results using our current approach. In the case of the concrete problem (Figure 14), our experimental network outperforms a standard Kaiming-initialized network when block diagonality is enforced. In the case of the two-spiral classification task (Figure 15), our network is able to learn an accurate decision boundary (left subfigure), whereas a standard network constrained to be block diagonal fails to learn (right subfigure). When dense weight matrices are trained, our networks are almost able to tie with the standard initializations, although convergence can take longer. These results suggest that our method might be quite powerful when its parameter budget is spent on greater width and depth of convex function blocks (as in Figure 8) instead of filling in dense matrices.
@@
A.11 Larger Image Classification Examples
We include very preliminary demonstrations of how our method can increase performance on large-scale image classification tasks. A nice feature of our method is that it can be used anywhere dense layers appear in networks, which enables its use in CNN architectures. Nonetheless, combining our construction with convolutions or other layers is a use case far outside of the scope of some of the toy problems where we observe orders of magnitude of improvement, and the output of each subnetwork might not learn to be as nicely behaved as in the tightly controlled experimental setting of 1D convex training data. Thus, we stress that while our initial results are positive, and we can reasonably conjecture why, future work on extending our method to arbitrary ReLU networks would provide greater rigor and likely far greater performance gains, as seen with our lower-dimensional function approximation examples that realize the full benefit of our mathematically principled approach.
We evaluated two standard benchmarks: CIFAR-10 and ImageNet. In the case of CIFAR-10, we trained a small network of three convolutions followed by two fully connected layers. To apply our technique, we inserted our custom networks in between the two fully connected layers to act as activations, and reduced the number of neurons slightly to compensate for the extra parameters added. In the case of ImageNet, we used PyTorch’s pretrained VGG16 model for its convolution weights (which were frozen during training). We modify the linear layers at the end of the network in the same manner as we did for the CIFAR-10 experiment, and compare against re-learning the fully connected weights with Kaiming initialization. We are thus able to evaluate the effect of our network (on the linear layers) versus Kaiming initialization, for both CIFAR-10 and ImageNet. Note that in both experiments, we switch parameterizations in our method halfway through training.
Figure 16: results for CIFAR-10 and ImageNet experiments. Our (“experimental”) training and validations generally remain below their Kaiming-initialized (“ordinary”) counterparts. Interestingly, the results on ImageNet are much stronger (note the color labeling is the same for both images, our train and validation losses are the bottom two in ImageNet, whereas in Cifar-10 the two training losses are the lowest). This may be because the ImageNet experiments re-used the convolutional layers from VGG, rather than learning them on its own. It could also be in part due to larger layer sizes, since most activation functions are universal with infinite width, any detrimental effect of constraining them into convex shapes may be softened. In our ImageNet experiment we trained both networks for 2 epochs. Our method required a lower learning rate ($0.00001$) than used in previous experiments. At epoch 2, our training and validation losses were 0.01761 and 0.01657, compared to 0.02518 and 0.02269 for the standard networks (a 30% and 27% reduction, respectively).
These two examples show that even for large-scale image classification tasks—far away from the analytic 1D function tests considered in the main body—our methodology still has the potential to yield noticeable benefits, highlighting the promise of extending the present work to more general network architectures as a future research direction.
Potential Broader Impact
This paper presents work whose goal is to enable more efficient neural networks. Although the present work is largely theoretical, future advances in this line of research could enable the use of much smaller networks in many practical applications, which could substantially mitigate the rapidly growing issue of energy usage in large learning systems.

### v7

@@ B.1 Learning Rates
Figure 14: Learning $y=x^{3}$ at various learning rates for 1000 epochs. The figure shows the log losses of the best model of 30 for networks trained according to the methods in this paper versus random initialization and gradient descent (“Default Network”). For high learning rates, neither learns because the steps are too large. For low learning rates (below $10^{-5}$), the steps are too small, and our networks are likely deriving an implicit advantage from being forcibly initialized to a convex function. Both methods are able to converge for learning rates in between $10^{-4}$ and $10^{-2}$; one could run for more epochs to see a similar advantage of our method for smaller learning rates.
@@
B.2 Real-World Data and Classification Problems
@@
Nonetheless, while we believe there is much room for future work to improve higher-dimensional results, Figures 15 and 16 already show impressive results using our current approach. In the case of the concrete problem (Figure 15), our experimental network outperforms a standard Kaiming-initialized network when block diagonality is enforced. In the case of the two-spiral classification task (Figure 16), our network is able to learn an accurate decision boundary (left subfigure), whereas a standard network constrained to be block diagonal fails to learn (right subfigure). When dense weight matrices are trained, our networks are almost able to tie with the standard initializations, although convergence can take longer. These results suggest that our method might be quite powerful when its parameter budget is spent on greater width and depth of convex function blocks (as in Figure 9) instead of filling in dense matrices.
@@
B.3 Image Classification on CIFAR-10
Figure 17: Results for CIFAR-10 experiments. Our (“experimental”) training and validations are generally comparable to their Kaiming-initialized (“ordinary”) counterparts. The small architectures are trained with Adam at a learning rate of $0.001$, while the VGG experiments are trained with SGD, following a learning rate schedule with a 10x drop at epochs 80 and 120. The ordinary network uses an initial learning rate of 0.1, and the experimental network uses 0.01. No parameterization switches occur with our method.
A nice feature of our method is that it can be used anywhere dense layers appear in networks, which enables its use in CNN architectures. Here we demonstrate its application for image classification on CIFAR-10 (Krizhevsky et al., 2009). Our networks are inserted into the dense layers of VGG-16, as well as a small-scale CNN architecture of 3 convolutions (channel depths of $3\rightarrow 32\rightarrow 64\rightarrow 128$ with $3\times 3$ kernels) followed by a single hidden layer dense classifier of 256 neurons, whose width was reduced slightly to 247 to pay for the extra parameters when our method was used. The results we get are comparable between the two approaches (similar to the ImageNet experiments in Section 6). Again we notice that our method requires a lower learning rate for stability, and that switching parameterizations does not significantly impact accuracy.
Appendix C Source Code Release
An implementation of our technique can be found at https://github.com/NREL/triangle_net.